Fuel “X” is one of the fuels in the table above. It has a heat of combustion value of 3329 kJmol-1.
Combustion of 0.44g of this fuel causes an increase in the temperature of 200.0g of water of 19.9oC.
Assuming negligible heat loss to the environment outside of the water, which of the following identifies fuel “X”?
1
Expert's answer
2020-07-22T05:50:05-0400

heat of combustion value of 3329 kJmol-1


Let Moles of fuel = ( 0.44 / M ) MOLE .

WHERE M IS MOLAR MASS OF FUEL .


So ., Total Energy = 3329 * 1000 * (0.44 / M .) J .


This Enegy Utilise By The Water = ( 200 ) * ( 4.2 ) * 19.9 J = 16716 j .


Total Energy = This Enegy Utilise By The Water .


3329 * 1000 * (0.44 / M .) J = 16716 J .


So , M = ( 3329 * 1000 * 0.44 ) / 16716


So , M = 87.62 g / mol .


Molar Mass of Gas = 87.6 g / mol.